OpenStudy (anonymous):

Solve by Factoring (Solve for SOLUTIONS): X^3+5X^2+6X=0

OpenStudy (jdoe0001):

\(\large x^3+5x^2+6x=0\implies \begin{array}{cccllll} x(x^2&+5x&+6)&=0\\ &\uparrow &\uparrow \\ &3+2&3\cdot 2 \end{array}\)

OpenStudy (anonymous):

Would it be X(X+2) (X+3)? I dont know i may be confused

OpenStudy (jdoe0001):

well.... do you know how to factor quadratics?

OpenStudy (anonymous):

like -B over (b)^2 - 4(a)(c)

OpenStudy (anonymous):

X^3+5X^2+6X=0 x(x^2+5x+6)= 0 x ( (x + 2) (x + 3) ) = 0 so x = 0, or x = -2, or x = -3

OpenStudy (anonymous):

you get it?

OpenStudy (anonymous):

Yes Awesome! I have 1 more problem. I;; start a new question so you can get another medal

OpenStudy (anonymous):

ok